OpenStudy (anonymous):

I have the answer can someone just check me please. Which of the following is f '(x) for f(x)=8 square root x

OpenStudy (whpalmer4):

\[f(x) = 8\sqrt{x} = 8x^{1/2}\] What's your answer?

OpenStudy (anonymous):

I got 4x^(1/2)

OpenStudy (whpalmer4):

Remember what I said earlier?

OpenStudy (anonymous):

would it be instead 4x^(-1/2)

OpenStudy (anonymous):

no no no I mean just figured it out I think -4x^(1/2)

OpenStudy (whpalmer4):

\[\frac{d}{du}[u^n] = nu^{n-1}\]

OpenStudy (whpalmer4):

Take your existing expression, multiply it by the value of the exponent, and subtract one from the exponent.

OpenStudy (anonymous):

right so don't you get 4x^(-1/2)

OpenStudy (whpalmer4):

Yes. But you changed your mind immediately...

OpenStudy (anonymous):

because I was still doing the math and I forgot what you said agian as I was doing it and added the 1 again instead of subtracting it.

OpenStudy (whpalmer4):

Here's another one for you: \[f(x) = x^{2/3}\]

OpenStudy (anonymous):

x^(-1/3)

OpenStudy (anonymous):

is that right

OpenStudy (whpalmer4):

"Take your existing expression, multiply it by the value of the exponent, and subtract one from the exponent." Did you do all of that correctly?

OpenStudy (anonymous):

yes and I got x^(1/3)

OpenStudy (whpalmer4):

Really. 1 * 2/3 = 1?

OpenStudy (anonymous):

oh so then (2/3)x^(1/3)

OpenStudy (whpalmer4):

Better, but still incorrect. 2/3 - 1 = 1/3?

OpenStudy (anonymous):

sorry I forgot the (-1/3)

OpenStudy (whpalmer4):

So, final, correct answer is?

OpenStudy (anonymous):

(2/3)x^(-1/3)

OpenStudy (whpalmer4):

and written as a fraction with positive exponents that would be?

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }x ^{-1/3}\]

OpenStudy (whpalmer4):

Is "-1/3" a positive exponent?

OpenStudy (anonymous):

how do you make it positive exponent?

OpenStudy (whpalmer4):

\[a^{-n} = \frac{1}{a^n}\]

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 }x ^{\frac{ 1 }{ 1/3 }}\]

OpenStudy (whpalmer4):

No. \[\frac{2}{3}x^{-1/3} = \frac{2}{3}*x^{-1/3} = \frac{2}{3}*\frac{1}{x^{1/3}} = \frac{2}{3x^{1/3}}\]

OpenStudy (anonymous):

oh ok.

OpenStudy (whpalmer4):

To convert an exponential from positive to negative or vice versa, change the sign of the exponent and take the reciprocal.

OpenStudy (anonymous):

ok thanks

OpenStudy (whpalmer4):

The answer might also be written as \[\frac{2}{3\sqrt[3]{x}}\] because \[x^{1/n} = \sqrt[n]{x}\]